The platter of the hard drive of a computer rotates at 7600 rpm (rpm = revolutions per minute = rev/min).What is the angular velocity (rad/s) of the platter?If the reading head of the drive is located 3.00 cm from the rotation axis, what is the linear speed of the point on the platter just below it?If a single bit requires 0.50 μm of length along the direction of motion, how many bits per second can the writing head write when it is 3.00 cm from the axis?
1
Expert's answer
2020-04-20T10:44:10-0400

a)


"f=\frac{7600}{60}\frac{rev}{s}"

"\omega=2\pi\frac{7600}{60}=796\frac{rad}{s}"

b)


"v = rω = (0.03)(754 ) = 22.6 \frac{rad}{s}"

c)  For each bit


"s = 0.5\cdot10^{-6} m"

"\frac{ 22.6 }{ 0.5\cdot10^{-6} }= 45\cdot10^{6} \frac{bit }{ s }"